# Primary gravitational waves at high frequencies I: Origin of suppression in the power spectrum

### v1

@@ 7.2 Outlook
There are a few directions in which the calculations we have carried out need to be extended. To begin with, we need to extend our analysis for transitions from inflation to radiation domination that are even more smooth. In fact, we have to work with $U(\eta)$ that are infinitely continuous. However, it is not easy to obtain solutions for the scalar factor $a(\eta)$ and the rescaled mode functions $\mu_{k}(\eta)$ in such situations. We may have to consider some approximations to determine the Bogoliubov coefficients $(\alpha_{k},\beta_{k})$ in these situations. Secondly, we need to introduce a phase of reheating between (inflation and the epoch of radiation domination) and evolve the PGWs across the phase of reheating as well as the epochs of radiation and matter domination, and evaluate the regularized PS of PGWs today. These are relatively straightforward exercises and can be implemented easily. Thirdly, we need to smooth the different transitions from the end of inflation until today and determine the corresponding imprints on the PS of PGWs. It seems clear that the suppression in the PS of PGWs over $k\gtrsim k_{\mathrm{e}}$ arises due to the smoothing of the transition from inflation to the epoch that succeeds it. The smoothing of the later transitions can be expected to lead to some features around wave numbers that re-enter the Hubble radius at the time of these transitions, say, $k_{\mathrm{re}}$, corresponding to the transition from the phase of reheating to the epoch of radiation domination, and $k_{\mathrm{eq}}$ corresponding to the time of transition from the epoch of radiation domination to that of matter domination. Fourthly, we have seen that the smoother transition from inflation to the epoch that follows will lead to a suppression in the regularized PS of PGWs over the wave numbers $k\gtrsim k_{\mathrm{e}}$. In the simplest of scenarios, a smooth transition from inflation to reheating and eventually to the epoch of radiation domination is achieved by coupling the inflation to radiation with the help of a constant coupling parameter. The manner in which the regularized PS of PGWs will be suppressed at high wave numbers can be expected to depend on the details of the transition. As is well known, the nature of the epoch of reheating will depend on the behavior of the inflationary potential near the minimum. It seems worthwhile to systematically examine the relation between the form of the suppression in the regularized PS and the behavior of the inflationary potential near its minimum and the coupling of the inflation to radiation. We are currently working on these problems.
@@ Appendix A Equation of motion of PGWs
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\tilde{h}_{\mu\nu}+j_{\mu\nu},$ (A.1)
where $\bar{g}_{\mu\nu}$ is the background metric, $\tilde{h}_{\mu\nu}$ is the first order perturbation in the metric, and $j_{\mu\nu}$ refers to the second order perturbations in the metric, i.e. it is of $\mathcal{O}(\tilde{h}^{2})$. Using this perturbative expansion, we can express the Ricci tensor as follows:
$R_{\mu\nu}=\bar{R}_{\mu\nu}+R_{\mu\nu}^{(1)}(\bar{g},\tilde{h})+R_{\mu\nu}^{(2)}(\bar{g},\tilde{h}^{2})+R_{\mu\nu}^{(2)}(\bar{g},j),$ (A.2)
where the first term is the Ricci tensor associated with the background, the second and third terms are linear and quadratic in $\tilde{h}_{\mu\nu}$, respectively, and the last term, as we have indicated, is the contribution from $j_{\mu\nu}$.
@@
where $\bar{p}$ is the background pressure that is driving the expansion. Note that, according to the zeroth order Friedmann equation,

### v2

@@ 7.2 Outlook
There are a few directions in which the calculations we have carried out need to be extended. To begin with, we need to extend our analysis for transitions from inflation to radiation domination that are even more smooth. In fact, we have to work with $U(\eta)$ that are infinitely continuous. However, it is not easy to obtain solutions for the scale factor $a(\eta)$ and the rescaled mode functions $\mu_{k}(\eta)$ in such situations. We may have to consider some approximations to determine the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ in these situations. Secondly, we need to introduce a phase of reheating (between inflation and the epoch of radiation domination) and evolve the PGWs across the period of reheating as well as the epochs of radiation and matter domination, and evaluate the regularized PS of PGWs today. These are relatively straightforward exercises and can be implemented easily. Thirdly, we need to smooth the different transitions from the end of inflation until today and determine the corresponding imprints on the PS of PGWs. It seems clear that the suppression in the PS of PGWs over $k\gtrsim k_{\mathrm{e}}$ arises due to the smoothing of the transition from inflation to the epoch that succeeds it. The smoothing of the later transitions can be expected to lead to some features around wave numbers that re-enter the Hubble radius at the time of these transitions, say, $k_{\mathrm{re}}$, corresponding to the transition from the phase of reheating to the epoch of radiation domination, and $k_{\mathrm{eq}}$ corresponding to the time of transition from the epoch of radiation domination to that of matter domination [108]. Fourthly, we have seen that the smoother transition from inflation to the epoch that follows will lead to a suppression in the regularized PS of PGWs over the wave numbers $k\gtrsim k_{\mathrm{e}}$. In the simplest of scenarios, a smooth transition from inflation to reheating and eventually to the epoch of radiation domination is achieved by coupling the inflaton to radiation with the help of a constant coupling parameter. The manner in which the regularized PS of PGWs will be suppressed at high wave numbers can be expected to depend on the details of the transition. As is well known, the nature of the epoch of reheating will depend on the behavior of the inflationary potential near the minimum. It seems worthwhile to systematically examine the relation between the form of the suppression in the regularized PS and the behavior of the inflationary potential near its minimum and the coupling of the inflaton to radiation. We are currently working on these problems.
@@ Appendix A Equation of motion of PGWs
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\tilde{h}_{\mu\nu},$ (A.1)
where $\bar{g}_{\mu\nu}$ is the background metric and $\tilde{h}_{\mu\nu}$ is the perturbation in the metric at the first order. Using this perturbative expansion, we can express the Ricci tensor as follows:
[…]
where the first term is the Ricci tensor associated with the background, while the second term is linear in $\tilde{h}_{\mu\nu}$.
@@
where $\bar{p}$ is the background pressure that is driving the expansion. Note that, according to the zeroth order Friedmann equation